OpenStudy (anonymous):

A straight line inserted at the origin terminates at the point (6, 7) as it sweeps out an angle θ in standard position. Which of the following corresponds to the evaluation of secθ;

OpenStudy (anonymous):

\[1.\frac{ \sqrt{85} }{ 6} 2.\frac{ 6\sqrt{85} }{ 85 } 3.\frac{ 7 }{ 6 } 4.\frac{ 6 }{ 13 }\]

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

The graph shows Susan’s distance from her office, in yards, as she runs errands, over time, in minutes. How much time did Susan spend moving toward her office?

OpenStudy (anonymous):

y would you put that

OpenStudy (anonymous):

dont post a question in my question

OpenStudy (anonymous):

The graph shows Susan’s distance from her office, in yards, as she runs errands, over time, in minutes. How much time did Susan spend moving toward her office?

OpenStudy (anonymous):

stop

OpenStudy (anonymous):

The graph shows Susan’s distance from her office, in yards, as she runs errands, over time, in minutes. How much time did Susan spend moving toward her office? min The graph has distance in yards on the y-axis and time in minutes on the x-axis. The graph has six sections. Section 1 is a line from point 0, 0 to point 6, 140. Section 2 is a line from point 6, 140 to point 10, 140. Section 3 is a line from point 10, 140 to point 12, 90. Section 4 is a line from point 12, 90 to point 16, 70. Section 5 is a line from point 16, 70 to point 22, 70. Section 6 is a line from point 22, 70 to point 26, 0.

OpenStudy (anonymous):

for real stop

OpenStudy (anonymous):

What is it

OpenStudy (anonymous):

i am not answering your question you posted it in my question make your own post

OpenStudy (anonymous):

how

OpenStudy (anonymous):

....just go away

OpenStudy (anonymous):

:I